Question A: Logarithmic Equations A1. (x 1)2 2. How many real solutions does it have? Solve the equation logz(x
Answer #1

A1. We have log3 (x-1)2 = 2

Therefore, (x-1)2 = 32 = 9 so that x-1 = -3 or, 3.

Then x = 1-3 = -2 or, 1+3 = 4.

Thus, there are 2 solutions, i.e. x = -2 or, 4.